OpenStudy (anonymous):

What operation is necessary to solve the equation that represents the statement: 2.4 is 8 multiplied by what number?

OpenStudy (anonymous):

What is the opposite of multiplication?

OpenStudy (anonymous):

division

OpenStudy (anonymous):

Represent this as an equation. 2.4 * Y = 8 where Y is the number that you want to find. Can you solve this equation?

OpenStudy (anonymous):

8 * 2.4= 0.3

OpenStudy (anonymous):

When you are solving an equation, you are trying to find a variable you do not know, in this case Y 2.4 * Y = 8 So you need to make Y the subject of the equation, in english this means that you must have just Y on one side of the equation and all the other stuff on the other side of the equation. So you need to get rid of 2.4* bit. How would you do that? Well you apply the opposite of multiplication which is division. And you apply it to both sides. So you divide both sides of the equation by 2.4 2.4 * Y = 8 2.4 * Y / 2.4= 8 / 2.4 Y = 8 / 2.4

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thnx